# Visual Context Window Extension: A New Perspective for Long Video Understanding

### v1

@@ 3.2 Progressive Pooling
Table 1: Performance evaluation on VideoMME (Fu et al., 2024) benchmark.
@@ 4.3 Quatitative Results
Table 3: Performance evaluation on LongVideoBench (Wu et al., 2024) benchmark. 1 fps* ($\leqslant$512) indicates that for videos longer than 60 seconds, Frame Rate Sampling (N fps) is used, which samples $N$ frames per second. The remaining frames are uniformly sampled to a total of 512 frames.
@@
1 fps* ($\leqslant$ 512) 70.4 73.3 57.8 52.8 59.4
Results on VideoMME Table 1 presents the results on the VideoMME benchmark. Compared to the baseline model, LLaVA-OneVision, our method shows consistent improvements across all intervals for short, medium, and long videos. Notably, for long videos, the accuracy improved by 2.3%. In comparison to the latest long video understanding models, our approach continues to achieve optimal performance. For instance, compared to LongVILA-8B, which was pre-trained on long video-text pairs, our method demonstrates an improvement of 9.9%. Crucially, our method achieves these gains without requiring any pre-training or fine-tuning on long video-text pairs.
Results on MLVU and LongVideoBench MLVU and LongVideoBench are two benchmarks specifically designed to evaluate long video understanding tasks. Table 2 presents the results on MLVU, where our method significantly outperforms all comparison models, even surpassing GPT-4o. Table 3 provides the results on LongVideoBench, where test samples are categorized into various duration intervals to highlight different models’ performance in long video comprehension. While our method shows a slight performance drop in the intervals (8, 15] and (15, 60] when sampling 512 frames compared to the baseline LLaVA-OneVision, it achieves notable improvements of 3.9% and 4.0% in the (180, 600] and (900, 3600] intervals, respectively. This performance drop in shorter intervals can be attributed to the fact that dense frame sampling results in excessively long input sequences for shorter videos, which leads to attention distraction and degrades model performance. When frames are sampled at 1 frame per second (fps), the model’s performance improves significantly in both the (8, 15] and (15, 60] intervals. Overall, compared to other models, our approach consistently demonstrates superior performance.
@@ 4.4 Ablation Studies
Ours (Tuning-free) w/o progressive pooling 256 71.5 59.1 52.2 61.0

### v2

@@ 3.2 Progressive Pooling
Table 1: Performance evaluation on VideoMME (Fu et al., 2024) benchmark. * indicates the results of reproduction.
@@ 4.3 Quatitative Results
Table 3: Performance evaluation on LongVideoBench (Wu et al., 2024) benchmark.
@@
Results on VideoMME Table 1 presents the results on the VideoMME benchmark. Compared to the baseline model, LLaVA-OneVision, our method shows consistent improvements across all intervals for short, medium, and long videos. Notably, for long videos, the accuracy improved by 3.2%. In comparison to the latest long video understanding models, our approach continues to achieve optimal performance. For instance, compared to LongVILA-8B, which was pre-trained on long video-text pairs, our method demonstrates an improvement of 10.8%. Crucially, our method achieves these gains without requiring any pre-training or fine-tuning on long video-text pairs.
Results on MLVU and LongVideoBench MLVU and LongVideoBench are two benchmarks specifically designed to evaluate long video understanding tasks. Table 2 presents the results on MLVU, where our method significantly outperforms all comparison models, even surpassing GPT-4o. Table 3 provides the results on LongVideoBench, where test samples are categorized into various duration intervals to highlight different models’ performance in long video comprehension. Our method shows a slight performance drop in the intervals (8, 15] and (15, 60] when sampling 512 frames compared to the baseline LLaVA-OneVision. This performance drop in shorter intervals can be attributed to the fact that dense frame sampling results in excessively long input sequences for shorter videos, which leads to attention distraction and degrades model performance. Using different frame sampling strategies for videos of varying durations can alleviate this issue.
@@ 4.4 Ablation Studies
Ours (Tuning-free) w/o progressive pooling 256 71.6 59.1 52.2 61.0